# Towards Unified Multi-Modal Personalization: Large Vision-Language Models for Generative Recommendation and Beyond

### v1

@@ 2.1 Incorporate heterogeneous user history information
We then replace the user history with the user sentence, ensuring no loss of information, and incorporate a special token [CLS] at the beginning of the sequence. Notably, unlike previous personalized systems (Geng et al., 2022; Li et al., 2023) using only text or item IDs, in this study, we aim to handle multi-modal heterogeneous user sequence. This sequence including visual information isn’t immediately compatible with the language model.
@@ 4 Related Work
Multi-modal Personalization. The development of multi-modal personalization systems is crucial due to the prevalence of multi-modal information on the Web. The most common usage is to leverage multi-modal content as side information to assist recommendation decisions. For example, VBPR (He & McAuley, 2016b) and AMR (Tang et al., 2019) propose using visual features to support collaborative filtering computation. MMGCN (Wei et al., 2019) designs a multi-modal Graph Convolution network framework built upon the message-passing idea of graph neural networks, which can yield modal-specific representations of users and micro-videos to better capture user preferences. CausalRec (Qiu et al., 2021) develops a causal inference framework to effectively make use of the visual feature and in the meanwhile remove the visual bias. MMSSL (Wei et al., 2023b) proposes a modality-aware interactive structure learning paradigm via adversarial perturbations for data augmentation to generate robust representations. However, all the above methods only focus on single personalized recommendation tasks and fail to model the mixture of visual, textual, and ID data, arranged in arbitrary sequences.
@@
General-purpose Model. Language is an incredibly powerful tool that can effectively describe a wide range of concepts. With the advancements in modern language models, language has evolved into a versatile interface that can be applied to various tasks and modalities. Several methods have emerged to connect different tasks and modalities within a unified language model. VL-T5 (Cho et al., 2021) and Pix2Seq (Chen et al., 2021) approach vision-language tasks and object detection by formulating them as sequence-to-sequence problems based on conditional language generation. To solve the text-to-image generation task, DALL-E (Ramesh et al., 2021), on the other hand, tackles the text-to-image generation task by converting images into latent tokens and employing a Transformer decoder to model both text and image tokens in an autoregressive manner. In a similar vein, OFA (Wang et al., 2022) has contributed to the development of comprehensive vision-language foundation models that excel in generating text from images. Moving towards personalization, P5 (Geng et al., 2022) has proposed a unified framework that integrates various recommendation tasks, while VIP5 (Geng et al., 2023) extends this framework to a multi-modal setting. However, these are limited to a narrow scope of tasks, and the dilemma of multi-task instruction optimization is not well addressed. They lack the required flexibility to effectively handle the demands of heterogeneous input and output requirements.
@@
Large language model for personalization. Personalization has been well-studied for information access problems. Recently, large language models (LLMs) have shown great performance in various personalization tasks. Some works explored personalization for fundamental language modeling with the openly available user data on Reddit (Welch et al., 2022), Facebook, and Twitter (Soni et al., 2022). (Soni et al., 2022) also explore applying a personalized language model for downstream tasks in stance classification and demographic inference. Similarly, other work has explored personalized sentiment prediction tasks on publicly available Yelp and IMDB data (Mireshghallah et al., 2022; Zhong et al., 2021). The large language model has demonstrated remarkable success in learning unified representations and has the potential to enhance the personalization of recommendation systems (Geng et al., 2022; Li et al., 2023; Hou et al., 2022) as well. (Salemi et al., 2023) introduces the LaMP benchmark and highlights the importance of personalization for LLMs. However, most previous methods mainly focus on single personalized recommendation tasks and fail to model the multi-modal data.
@@ A.2 Implementation Details
[…]
@@ A.3 Robustness of Proposed Method to Missing Modality (RQ5)
[…]

### v2

@@ 2.1 Incorporate heterogeneous user history information
We then replace the user history with the user sentence, ensuring no loss of information, and incorporate a special token [CLS] at the beginning of the sequence. Notably, unlike previous personalized systems (Geng et al., 2022; Li et al., 2023b) using only text or item IDs, in this study, we aim to handle multi-modal heterogeneous user sequence. This sequence including visual information isn’t immediately compatible with the language model.
@@ 4 Related Work
Multi-modal Personalization. The development of multi-modal personalization systems is crucial due to the prevalence of multi-modal information on the Web. The most common usage is to leverage multi-modal content as side information to assist recommendation decisions. For example, VBPR (He & McAuley, 2016b) and AMR (Tang et al., 2019) propose using visual features to support collaborative filtering computation. MMGCN (Wei et al., 2019) designs a multi-modal Graph Convolution network framework, which can yield modal-specific representations of users and micro-videos to better capture user preferences. CausalRec (Qiu et al., 2021) develops a causal inference framework to effectively make use of the visual feature and in the meanwhile remove the visual bias. MMSSL (Wei et al., 2023b) proposes a modality-aware interactive structure learning paradigm via adversarial perturbations to generate robust representations. However, all the above methods only focus on single personalized recommendation tasks and fail to model the mixture of visual, textual, and ID data, arranged in arbitrary sequences.
@@
General-purpose Model. Language is an incredibly powerful tool that can effectively describe a wide range of concepts. With the advancements in modern language models, language has evolved into a versatile interface that can be applied to various tasks and modalities. Several methods have emerged to connect different tasks and modalities within a unified language model. VL-T5 (Cho et al., 2021) and Pix2Seq (Chen et al., 2021) approach vision-language tasks and object detection by formulating them as sequence-to-sequence problems based on conditional language generation. To solve the text-to-image generation task, DALL-E (Ramesh et al., 2021), on the other hand, tackles the text-to-image generation task by converting images into latent tokens and employing a Transformer decoder to model both text and image tokens in an autoregressive manner. In a similar vein, OFA (Wang et al., 2022) has contributed to the development of comprehensive vision-language foundation models that excel in generating text from images. Flamingo (Alayrac et al., 2022; Awadalla et al., 2023), BLIP (Li et al., 2023c) and Otter (Li et al., 2023a) further advance multi-modal foundation models, exploring robust modal alignment and instruction tuning mechanisms. Moving towards personalization, P5 (Geng et al., 2022) has proposed a unified framework that integrates various recommendation tasks, while VIP5 (Geng et al., 2023) extends this framework to a multi-modal setting. However, they fail to fully harness the potential of raw data and do not effectively capture the interactions among them. They lack the required flexibility to effectively handle the diverse input and output requirements inherent in multi-task learning.
@@
Large language model for personalization. Personalization has been well-studied for information access problems. Recently, large language models (LLMs) have shown great performance in various personalization tasks. Some works explored personalization for fundamental language modeling with the openly available user data on Reddit (Welch et al., 2022), Facebook, and Twitter (Soni et al., 2022). (Soni et al., 2022) also explore applying a personalized language model for downstream tasks in stance classification and demographic inference. Similarly, other work has explored personalized sentiment prediction tasks on publicly available Yelp and IMDB data (Mireshghallah et al., 2022; Zhong et al., 2021). The large language model has demonstrated remarkable success in learning unified representations and has the potential to enhance the personalization of recommendation systems (Geng et al., 2022; Li et al., 2023b; Hou et al., 2022) as well. (Salemi et al., 2023) introduces the LaMP benchmark and highlights the importance of personalization for LLMs. However, most previous methods mainly focus on the text and fail to model the multi-modal data.
@@ A.2 Implementation Details
We compare with the representative and state-of-the-art baselines of each task. Notably, P5 (Geng et al., 2022) and VIP5 (Geng et al., 2023) are unified models that can work on partial of tasks. For recommendation, we split the baselines into four categories - general recommenders: MF (Koren et al., 2009; Rendle et al., 2012), MACR (Wei et al., 2021), LightGCN (He et al., 2020), UltraGCN (Mao et al., 2021). Sequential recommenders: HGN (Ma et al., 2019), GRU4Rec (Jannach & Ludewig, 2017), SASRec (Kang & McAuley, 2018) and S3-Rec (Zhou et al., 2020). Generative recommenders: BERT4Rec (Sun et al., 2019), UniSRec (Hou et al., 2022), P5 (Geng et al., 2022) and VIP5 (Geng et al., 2023). Multi-modal recommenders: VBPR (He & McAuley, 2016b), CausalRec (Qiu et al., 2021), MMGCL (Yi et al., 2022) and MMSSL (Wei et al., 2023b). In addition the recommenders, we use MRG (Truong & Lauw, 2019) as baseline for explaination generation, SANCL Han et al. (2022) for preference prediction, S3-Rec (Zhou et al., 2020) and UniSRec (Hou et al., 2022) utilizing threshold-based method for multi-modal selection, BM25 (Robertson et al., 2009), HEM (Ai et al., 2017), ZAM (Ai et al., 2019), RTM (Bi et al., 2021). Note that general recommenders operate differently compared to other types and are included primarily to ensure a thorough and comprehensive comparison. It’s important to note that the other categories (e.g., generative recommenders) are our primary baselines for comparison. We leave the comparisons with multi-view learning methods (Fu et al., 2020) as our future work.
@@ A.3 Robustness of Proposed Method to Missing Modality (RQ5)
In Table 6, we present the robustness of UniMP in the face of noisy textual and visual inputs. Here, NT represents textual input in which 20% of the attribute information is absent, and NV denotes visual input where 20% of the images do not match their corresponding product. Both of these scenarios are reflective of real-world application conditions. It is evident from the table that UniMP is robust to such noisy input data, and it will effectively leverages the other modality to deliver accurate predictions with minimal performance loss. On the other hand, the performance of the baseline methods MMSSL and S3-Rec deteriorates considerably in these conditions.